Grade 12Physical Chemistry

A solution contains 16g of methanol and 90g of water the mole fraction of methanol in the solution is?

Roshan Nayar
7 Years agoGrade 12

2 Answers

Arun
7 Years ago
moles of solute = 16 g / 32 g / mole = 0.50 moles 
moles of solvent = 90.0 g / 18 g / mol = 5.0 moles  

total number of moles of solute + solvent = 5.50 moles  

mole fraction of the methanol = 0.50 / 5.50 = 0.1
Khimraj
7 Years ago
Methanol = CH4 O  or  CH3 OH
molar mass of methanol:  12+4+16=32
       =>   16 gram =  0.5 moles
90 grams of water =>    5 mole of water as molar mass of water is 18.
total number of moles in the solution = 5.5 
so molar fraction of methanol in the solution =  0.5 /5.5 = 1/11